OpenStudy (anonymous):

solve for s please ( -12s)-11=(-11s)+4

OpenStudy (anonymous):

@Parvathysubhash

OpenStudy (anonymous):

@apoorvk i hav gt a doubt on dis

OpenStudy (anonymous):

i men that...........there i edited the question.........@Parvathysubhash

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

no u ask @apoorvk

OpenStudy (anonymous):

@apoorvk hey..............i need somee help

OpenStudy (anonymous):

@Diyadiya @ParthKohli @rajathsbhat ..........help??

OpenStudy (diyadiya):

(-12s)-11=(-11s)+4 -12s -11 = -11s +4 First add 11s on both sides

OpenStudy (anonymous):

k

OpenStudy (anonymous):

-12s=+4=3 or -3 ??

OpenStudy (diyadiya):

-12s -11 +11s = -11s +4 +11s -s -11= 4 Now add 11 on both sides!

OpenStudy (diyadiya):

No, What you did is wrong! @barbiejazzybabii

OpenStudy (diyadiya):

You can only add like terms you can't add 11s & 11

OpenStudy (diyadiya):

So continue..

OpenStudy (anonymous):

so...........23s-11+4

OpenStudy (diyadiya):

What??

OpenStudy (diyadiya):

-12s -11 +11s = -11s +4 +11s -s -11= 4 did you understand this?

OpenStudy (apoorvk):

*facepalm*

OpenStudy (anonymous):

not really....no @apoorvk......wtf is that??

OpenStudy (diyadiya):

Okay, This is your question -12s -11 = -11s +4 Now add 11s on both sides can you do that?

OpenStudy (apoorvk):

this tf is a 'facepalm': http://blog.domonet.co.uk/wp-content/uploads/2010/04/facepalm.jpg

OpenStudy (apoorvk):

@okay @barbiejazzybabii - what is x + 2x?

OpenStudy (anonymous):

oh it means duhhhh! haha not everyone is as smart as u.......ya @Diyadiya

OpenStudy (apoorvk):

@barbiejazzybabii no offences there, take your own time :)

OpenStudy (apoorvk):

You have algebra1?

OpenStudy (anonymous):

@apoorvk ......2x? .............idk cuz theres only one number,,,what plus two times what?..........doesnt really make sense.........idk im going to be a junior..im in summer school with pre algebra STILL

OpenStudy (anonymous):

im very good at english and spelling and grammar.........math is just not my thing at all.

OpenStudy (diyadiya):

Ok x+2x =3x (its x is considered as 1x so 1x + 2x =3x) But, You can't add 1 + 2x so what is 2x -x = ?

OpenStudy (apoorvk):

x + 2x = x(1 + 2) = x ( 3) = 3x so it's 3x basically, do you understand how? (Don't fret, don't get unnerved, stay cool, and go through this whole post again if you need)

OpenStudy (anonymous):

k im starting to get it........the x serves as a one

OpenStudy (diyadiya):

\(x\) is same as \(1x\) or \( 1 \times x =1x =x \) :) so tell me what is 2x -x = ?

OpenStudy (anonymous):

1x?

OpenStudy (diyadiya):

Yes! that's right! :) 1 more example what is 3x - 4x +2 = ?

OpenStudy (anonymous):

7x+2=9x?

OpenStudy (diyadiya):

Nope! that's where you're going wrong you can't add constants & variables together here 2 is a constant and a number with variable such as x, y ,a ,b ,c etc are variables so 3x -4x +2 = -x + 2 understood? another example 7y - 2y +4 +3 = ?

OpenStudy (diyadiya):

any doubts? My explanations are not really good ^_^

OpenStudy (apoorvk):

Not good? yes. Actually, Brilliant! Wish I could learn the virtue of patience from you Diya.

OpenStudy (diyadiya):

ROFL!! Thanks anyways~

OpenStudy (apoorvk):

That was nothing to ROFL about. Am offended. =_=

OpenStudy (apoorvk):

xD

OpenStudy (diyadiya):

i forgot "No Offence" :D Btw @barbiejazzybabii What happened?

OpenStudy (apoorvk):

@barbiejazzybabii please don't faint! *mumbles imaginary prayers*. :P

OpenStudy (anonymous):

oh oops sorry

OpenStudy (diyadiya):

Please do read my above post ^_^

OpenStudy (anonymous):

ahh whatever ill just get it wrong its confusing

OpenStudy (diyadiya):

No, you won't be able to do the other problems if you don't understand this just remember you can't add constants and variables together like 2x +3 is equal 2x+3 (you can't add them as 5x ) but 2x + 3x = 5x so can you tell me what is 7x + 2x +9 = ?

OpenStudy (anonymous):

5x+9 -5x-5 4x

OpenStudy (anonymous):

x=4???

OpenStudy (diyadiya):

Nooo! i'm just telling you to add them you can't find x because you don't have anything on the other side of equal to sign! 7x + 2x =9x so 7x + 2x +9 = 9x + 9

OpenStudy (diyadiya):

just tell me what is 3x + 5x = ?

OpenStudy (anonymous):

8x

OpenStudy (anonymous):

ahh ok i gotta get this question done ive been on it for an hour

OpenStudy (anonymous):

sooooooooooooooo................. ill just guess hahah bye.

OpenStudy (apoorvk):

What is 2 apples + 4 oranges? :\

OpenStudy (anonymous):

6 dumbass

OpenStudy (anonymous):

oopss that was my bro sorry....but yeah 6 fruitss...why do you have to put it like that?? im not in kindergarten............andd im DONE with this question it took me over an hour sooooooooooooo..................

OpenStudy (apoorvk):

Great that you're done now!!! :) And LOLOLOL at your brother's post! :P